# Total Gluon Helicity Contribution to Proton Spin from Lattice QCD

### v1

@@ Abstract
We report a state-of-the-art lattice QCD calculation of the total gluon helicity contribution to proton spin, $\Delta G$. The calculation is done on ensembles at three different lattice spacings $a=\{0.08,0.09,0.11\}$ fm. By employing distillation + momentum smearing for proton external states, we extract the bare matrix elements of the topological current $K^{\mu}$ under the 5-HYP smeared Coulomb gauge fixing configurations. Furthermore, we apply a non-perturbative $\mathrm{RI/MOM}$ renormalization scheme augmented with the Cluster Decomposition Error Reduction (CDER) technique to determine the renormalization constants of $K^{\mu}$. The results obtained from different components $K^{t,i}$ (with $i$ being the direction of proton momentum or polarization) are consistent with Lorentz covariance within uncertainties. After extrapolating to the continuum limit, $\Delta G$ is found to be $\Delta G=0.231(17)^{\mathrm{sta.}}(33)^{\mathrm{sym.}}$ at the $\overline{\mathrm{MS}}$ scale ${\mu}^{2}=10\ \mathrm{GeV}^{2}$, which constitutes approximately $46(7)\%$ of the proton spin.
@@
Introduction: Since the European Muon Collaboration (EMC) measured the quark spin contribution to proton spin through deep inelastic scattering with polarized muon and proton 1988364, the origin of proton spin has remained one of the most profound puzzles in modern particle and nuclear physics Ji:2020ena. Our current understanding is that the quark spin contribution accounts for only $\sim 30\%$ of the proton spin deFlorian:2009vb; Nocera:2014gqa; COMPASS:2015mhb; Alexandrou:2017oeh; Lin:2018obj; Liang:2018pis. The remaining contributions come from quark orbital angular momentum (OAM), as well as gluon spin and OAM. Two popular spin sum rules have been proposed in the literature to understand the proton spin structure. One is the Jaffe-Manohar sum rule Jaffe:1989jz which offers a complete decomposition of the proton spin into quark and gluon spin and OAM. It has a clear partonic picture in the infinite momentum frame (IMF) of the proton and lightcone gauge. The other is the frame- and gauge-independent sum rule by Ji Ji:1996ek, where the quark and gluon total angular momentum can be connected to the moments of generalized parton distributions Muller:1994ses; Ji:1996ek.
@@
Among these contributions, the understanding of gluon spin has posed a significant challenge, mainly due to the absence of a local gauge-invariant gluon spin operator Ji:2012gc. Nevertheless, it can be measured in polarized scattering experiments through the spin-dependent gluon helicity distribution $\Delta g(x)$ as its lowest moment $\Delta G=\int dx\Delta g(x)$. Note that unlike in the case of the quark distribution, the lowest moment of the gluon helicity distribution is still nonlocal. Only in the lightcone gauge does it reduce to the matrix element of a local operator $\vec{E}\times\vec{A}$. Phenomenological global analyses of experimental data indicate that $\Delta G$ is approximately 0.2 at the $\overline{\mathrm{MS}}$ scale $\mu^{2}=10\ \mathrm{GeV}^{2}$, provided that contributions from the small-$x$ region ($x<0.05$) with large uncertainties are excluded deFlorian:2009vb; deFlorian:2014yva. On the other hand, recent theoretical developments Ji:2013fga; Hatta:2013gta; Ji:2014lra, particularly within the framework of Large-Momentum Effective Theory (LaMET) Ji:2013dva; Ji:2014gla; Ji:2020ect, have enabled a direct computation of $\Delta G$ using lattice QCD. In this approach, instead of computing in the IMF and lightcone gauge which are both inaccessible on the lattice, one computes the large-momentum nucleon matrix element of a static gluon spin operator, $\vec{E}\times\vec{A}$, in a physical gauge that leaves the transverse gauge field intact. Actually, there exists an infinite number of operators that can be used to extract $\Delta G$, all belonging to the same universality class Hatta:2013gta. The static gluon spin matrix element can then be matched to $\Delta G$ through a perturbative factorization or matching relation Ji:2013fga. Based on this, the $\rm\chi QCD$ collaboration conducted an exploratory calculation of $\Delta G$ by evaluating the nucleon matrix element of $\vec{E}\times\vec{A}$ in the Coulomb gauge, yielding a result of $0.251(47)(16)$ Yang:2016plb. However, the one-loop lattice perturbative renormalization employed in that study introduces a large correction, indicating the necessity of a proper nonperturbative renormalization – which was, at the time, very challenging to implement due to the Coulomb gauge fixing. Furthermore, recent studies Pang:2024sdl have revealed that the LaMET factorization for $\Delta G$ is incompatible with that for $\Delta g(x)$.
@@
To circumvent the difficulties mentioned above, two new proposals have been made in Ref. Pang:2024sdl. One is based on using the topological current, and the other employs appropriate static nonlocal gluon operators generating the gluon helicity distribution. Both of them have a trivial matching only, and therefore resolve the inconsistency previously encountered in the matching for $\Delta G$ and $\Delta g(x)$. Moreover, their nonperturbative renormalization can be easily implemented. This opens a new avenue for systematic studies of $\Delta G$ from lattice QCD.
@@
In this work, we present a state-of-the-art lattice calculation of the total gluon helicity contribution to proton spin, $\Delta G$, using the topological current approach proposed in Ref. Pang:2024sdl. The calculation is done on ensembles at three different lattice spacings $a=\{0.08,0.09,0.11\}$ fm, which allow us to perform a continuum extrapolation. Distillation HadronSpectrum:2009krc, momentum smearing Bali:2016lva, and cluster decomposition Liu:2017man techniques have been used to improve the signal-to-noise ratio of proton matrix elements and parton Green’s functions, respectively. The nonperturbative renormalization has been performed in the $\mathrm{RI/MOM}$ scheme, following the operator mixing pattern given in Ref. Pang:2024sdl. After converting to the $\overline{\rm MS}$ scheme, the result is then evolved to the scale $\mu^{2}=10\,{\rm GeV}^{2}$ using the renormalization group (RG) equation with three-loop anomalous dimensions.
@@
Theoretical Framework: Following Ref. Pang:2024sdl, the total gluon helicity $\Delta G$ can be extracted by computing the matrix element of the local topological current
@@
in a longitudinally polarized proton. More specifically, the proton matrix element of the Coulomb gauge fixed operator $K^{\mu}$ ($\mu=t,i$ with $i$ denoting the component along the direction of proton momentum or polarization) reduces to $\Delta G$ in the IMF, namely,
$\frac{\langle\mathrm{PS}_{\mathrm{Proj.i}}|\int d^{3}xK^{t/i}_{\mathrm{C.G.}}(x)|\mathrm{PS}_{\mathrm{Proj.i}}\rangle}{2S^{t/i}}|_{\mathrm{IMF}}=\Delta G,$ (2)
where $|\mathrm{PS}_{\mathrm{Proj.i}}\rangle$ denotes the proton external state with specific momentum and polarization direction $i$, while $S^{t/i}$ represents the corresponding polarization vector. The subscript $\mathrm{C.G.}$ indicates the Coulomb gauge fixing for the topological current.
As shown in Ref. Pang:2024sdl, the ultraviolet (UV) divergence of $K^{\mu}$ is gauge independent. Therefore, we can perform a nonperturbative renormalization of $K^{\mu}$ in the $\mathrm{RI/MOM}$ scheme MARTINELLI199581 where the renormalization factor is typically computed in the Landau gauge. $K^{\mu}$ mixes only with the axial-vector current $J^{\mu}_{5}=\bar{q}_{i}\gamma_{5}\gamma^{\mu}q_{i}$ even under RI/MOM renormalization Pang:2024sdl. The renormalization constants can be determined by the following conditions
$\begin{pmatrix}\langle ps|K^{\mu}|ps\rangle^{\mathrm{tree}}\\ \langle ps|J^{\mu}_{5}|ps\rangle^{\mathrm{tree}}\end{pmatrix}=\begin{pmatrix}Z_{11}^{\mathrm{RI}}&Z_{12}^{\mathrm{RI}}\\ Z_{21}^{\mathrm{RI}}&Z_{22}^{\mathrm{RI}}\end{pmatrix}\begin{pmatrix}\langle ps|K^{\mu}|ps\rangle^{\mathrm{lat.}}\\ \langle ps|J^{\mu}_{5}|ps\rangle^{\mathrm{lat.}}\end{pmatrix}$, (3)
where $|ps\rangle$ denotes either gluon ($g$) or quark ($q$) external states with specific momentum and polarization. In the discussion below, we introduce the abbreviations $\mathcal{O}^{\mathrm{tree/lat.},{g/q}}\equiv\langle{g,q}|\mathcal{O}|{g,q}\rangle^{\mathrm{tree/lat.}}$ to simplify the notation.
Note that $K^{\mathrm{tree},q}=J_{5}^{\mathrm{tree},g}=0$. Moreover, lattice data indicate that $J_{5}^{\mathrm{lat.},g}\approx 0$ within statistical uncertainties. Thus, the renormalization constants can be determined as:
@@
where $\mu_{\rm RI}$ denotes the renormalization scale in the RI/MOM scheme. With the above approximation, $Z_{22}^{\mathrm{RI}}$ reduces to the axial-vector current renormalization constant $Z_{A}^{\mathrm{RI}}$ in Ref. Hu:2023jet. The ratio $K^{\mathrm{lat.,q}}/J_{5}^{\mathrm{lat.,q}}$ appearing in the off-diagonal element $Z_{12}^{\mathrm{RI}}$ is of $\mathcal{O}(\alpha_{s})$. Thus, the $\mathrm{RI/MOM}$-renormalized gluon and quark helicities are given by:
@@
The RI/MOM renormalized results can then be converted to the $\overline{\mathrm{MS}}$ scheme via Pang:2024sdl
@@
where again the scale dependence has been suppressed for simplicity. In the present work, we have included the conversion factors up to three-loops. Their detailed expressions are given in the Supplemental Material supplemental.
Lattice Calculation: Our calculation is carried out on the $n_{f}=$2+1 flavor full QCD ensembles generated by CLQCD using the tadpole improved tree level Symanzik (TITLS) gauge action and the tadpole improved tree level Clover (TITLC) fermion action Hu:2023jet. The ensembles used for the simulation are described in Table 1. The values of $m_{\pi}$ of all the ensembles are chosen to ensure $m_{\pi}L\geq 3.7$, thereby efficiently suppressing the finite-volume effect. For more details, we refer the readers to the Supplemental Material supplemental.
@@
Table 1: Gauge coupling $\hat{\beta}=10/(g_{0}^{2}u_{0}^{4})$, lattice spacing $a$, dimensionless lattice size $\tilde{L}^{3}\times\tilde{T}$, corresponding pion mass $m_{\pi}$, and number of configurations $N_{\mathrm{cfg}}$ used for the simulation.
@@
For proton external states at low momenta ($p\leq 0.5\ \mathrm{GeV}$), we implement the distillation smearing technique HadronSpectrum:2009krc, while for higher momentum states we employ an approach combining distillation with momentum smearing Bali:2016lva. We insert the topological current operator into the two-point function of a proton polarized along the $i$th-direction (denoted as $\mathrm{Proj.i}$) to obtain the disconnected three-point function
$C_{3}(t_{f},t_{i},t_{s})\equiv\langle\mathrm{PS}_{\mathrm{Proj.i}}(t_{f})|K^{t/i}_{\mathrm{C.G.}}(t_{i})|\mathrm{PS}_{\mathrm{Proj.i}}(t_{s})\rangle,$ (7)
where $t_{f}$, $t_{i}$, and $t_{s}$ correspond to sink, current insertion, and source time-slices, respectively. The bare matrix elements (BMEs) of the topological current are extracted through the following ratio:
@@
with $C_{2}(t_{f},t_{s})\equiv\langle\mathrm{PS}_{\mathrm{UnProj.}}(t_{f})|\mathrm{PS}_{\mathrm{UnProj.}}(t_{s})\rangle$ representing the unpolarized proton two-point correlation function. To optimize signal quality, we perform a summation over all source time-slices: $R_{K^{\mu}}(t_{f},t_{i})=\sum_{t_{s}}R_{K^{\mu}}(t_{f},t_{i},t_{s})$, which is then fitted to the form
$R_{K^{\mu}}(t_{f},t_{i})=\langle K^{\mu}\rangle+c_{1}e^{-\Delta E(t_{f}-t_{i})}+c_{2}e^{-\Delta Et_{i}},$ (9)
where $\langle K^{\mu}\rangle$ serves as an abbreviation for $\langle\mathrm{PS}_{\mathrm{Proj.i}}|K^{\mu}|\mathrm{PS}_{\mathrm{Proj.i}}\rangle/(2E)$, and $c_{1},c_{2},\Delta E$ are fitting parameters. For illustrative purposes, we show in FIG. 1 the fitted ground state matrix element for both $K^{i}$ and $K^{t}$ at $p=0.98$ GeV. The fitting results for other momenta can be found in the Supplemental Material supplemental.
@@
Figure 1: The ratio $R_{K^{\mu}}(t_{f},t_{i})$ at $t_{f}\in[0.63,1.16]\ \mathrm{fm}$ and the fitted BME $\langle K^{\mu}\rangle$ (the gray error band) for proton external states with p=0.98GeVp=0.98\ $\mathrm${GeV} under ensemble C24P29, utilizing the fitting procedure described in (9).
@@
$S_{g}^{\mu\nu}(p)=\mathrm{CV}\langle\mathrm{Tr}[A^{\mu}(p)A^{\nu}(-p)]\rangle,$ (11)
@@
where $\mathrm{CV}=2/[(N^{2}_{c}-1)\tilde{V}$], with $N_{c}=3$ representing the number of colors and $\tilde{V}=\tilde{L}^{3}\times\tilde{T}$ denoting the dimensionless lattice volume. Due to rotational symmetry breaking on the lattice, we can define the gluon field renormalization constants through two distinct prescriptions:
where $\hat{p}_{\mu}=\frac{2}{a}\sin{\frac{p_{\mu}a}{2}}$ represents the lattice momentum, and $\mu_{\rm RI}^{2}=p^{2}$ denotes the RI/MOM renormalization scale.
@@
$Z_{11}^{\mathrm{RI}}$ in Eq.(Total Gluon Helicity Contribution to Proton Spin from Lattice QCD) then takes the following form
@@
where the external momenta are chosen to satisfy $p_{1}=p_{2}=p$. The Green’s function of $K^{\mathrm{tree,g}}$ reads $i\epsilon^{\rho\mu\sigma\nu}\hat{p}_{\sigma}$ in our normalization convention. In addition, we hide the indicator of the bare Green’s function $G_{K}$ and bare gluon propagator $S_{g}$ in (13).
The inverse of $Z_{11}^{\rm RI}$ then reads,
$Z_{11}^{-1,\mathrm{RI}}(\mu_{\rm RI}^{2})$ $=$ $\sum_{\mu,\rho,\nu,\sigma}\frac{2\mathrm{CV}\hat{p}^{2}G_{K}^{\mu\rho\nu}(p,p)}{i\epsilon^{\rho\mu\sigma\nu}\hat{p}_{\sigma}(S_{g}^{\mu\mu}+S_{g}^{\nu\nu})(p^{2})}|_{\mu_{\rm RI}^{2}=p^{2}}$ (14)
$=$ $\sum_{\mu,\rho,\nu,\sigma}\frac{\mathrm{CV}\hat{p}^{4}G_{K}^{\mu\rho\nu}(p,p)}{i\epsilon^{\rho\mu\sigma\nu}\hat{p}_{\sigma}\hat{p}_{\mu}\hat{p}_{\nu}S_{g}^{\mu\nu}(p^{2})}|_{\mu_{\rm RI}^{2}=p^{2}}.$
@@
The summation over indices $\mu,\nu,\rho,\sigma$ in conjunction with the Levi-Civita antisymmetric tensor yields a total of 12 distinct configurations. To enhance the signal quality of the pure gluon Green’s function $G_{K}(p,p)$ in (14), we employ the cluster decomposition error reduction (CDER) technique Liu:2017man, which requires large volume ensembles. Due to the limited volume of C24P29, we have chosen to compute the renormalization constants on a larger volume ensemble, C48P23, which shares the same lattice spacing. In contrast, for the other two ensembles E32P29 and F32P30, the CDER can be directly applied. For more details on the CDER implementation, we refer the readers to the Supplemental Material supplemental.
@@
The conversion factors from the RI/MOM scheme to the $\overline{\rm MS}$ scheme are evaluated by choosing $\mu/\mu_{\rm RI}=e^{-{5}/{6}}$ Brodsky:1982gc in the fixed-order expression up to three-loops, and then RG-evolved to $\mu^{2}=10\,{\rm GeV}^{2}$. Subsequently, the discretization effects of ${\cal O}(a^{2}p^{2})$ are removed following the strategy proposed in Ref. Yang:2018nqn. Note that here the extrapolation is conducted by varying $p^{2}$ (and thus $a^{2}p^{2}$) and determining the value at $a^{2}p^{2}=0$. To simultaneously ensure the reliability of RI/MOM renormalization and an acceptable signal-to-noise ratio after applying the CDER, we choose the data in the momentum interval $\mu_{\mathrm{RI}}\in[3.0,5.0]\ \mathrm{GeV}$ for the extrapolation, which correspond to $a^{2}p^{2}\in[2.6,7.1]$ under ensemble C48P23. The extrapolated result for the inverse of $\bar{Z}_{11}^{\mathrm{lat}}\equiv R_{11}Z_{11}^{\mathrm{RI}}$ is shown in Fig. 2. Similar analyses have also been performed for the other two ensembles with different lattice spacings.
@@
Regarding the mixing with quark spin $\Delta\Sigma$, it turns out to be an order of magnitude less than $\bar{Z}_{11}^{\mathrm{lat}}$, especially at large $a^{2}p^{2}$ (See the Supplemental Material supplemental for more details). These properties offer a significant simplification in our numerical analysis, leading us to

### v2

# Total Gluon Helicity Contribution to the Proton Spin from Lattice QCD

@@ Abstract
We report a state-of-the-art lattice QCD calculation of the total gluon helicity contribution to the proton spin, $\Delta G$. The calculation is done on ensembles with three different lattice spacings $a=\{0.08,0.09,0.11\}$ fm. By employing distillation and momentum smearing for proton external states, we extract the bare matrix elements of the topological current $K^{\mu}$ using 5-HYP smeared Coulomb gauge fixing configurations. Furthermore, we apply a non-perturbative $\mathrm{RI/MOM}$ renormalization scheme augmented by the Cluster Decomposition Error Reduction (CDER) technique to determine the renormalization constants of $K^{\mu}$. The results obtained from different components $K^{t,i}$ (with $i$ being the direction of proton momentum or polarization) are consistent with Lorentz covariance within uncertainties. After extrapolating to the continuum limit, $\Delta G$ is found to be $\Delta G=0.231(17)^{\mathrm{sta.}}(44)^{\mathrm{sym.}}$ at the $\overline{\mathrm{MS}}$ scale ${\mu}^{2}=10\ \mathrm{GeV}^{2}$, which constitutes approximately $46(9)\%$ of the proton spin.
@@
Introduction: Since the European Muon Collaboration (EMC) measured the quark spin contribution to the proton spin using deep inelastic scattering of polarized muons and protons et al (1988), the origin of proton spin has remained one of the most profound puzzles in modern particle and nuclear physics Ji et al. (2021b). Our current understanding is that the quark spin contribution accounts for only $\sim 30\%$ of the proton spin de Florian et al. (2009); Nocera et al. (2014); Adolph and others (2016); Alexandrou et al. (2017); Lin et al. (2018); Liang et al. (2018). The remaining contributions come from quark orbital angular momentum (OAM), as well as gluon spin and OAM. Two popular spin sum rules have been proposed to help understanding the proton spin structure. One is the Jaffe-Manohar sum rule Jaffe and Manohar (1990) which offers a complete decomposition of the proton spin into quark and gluon spin and OAM. It has a clear partonic picture in the infinite momentum frame (IMF) of the proton and lightcone gauge. The other is the frame- and gauge-independent sum rule by Ji Ji (1997), where the quark and gluon total angular momentum can be connected to the moments of generalized parton distributions Müller et al. (1994); Ji (1997).
@@
Among these contributions, the understanding of gluon spin has posed a significant challenge, mainly due to the absence of a local gauge-invariant gluon spin operator Ji et al. (2012). Nevertheless, it can be deduced in polarized scattering experiments from the spin-dependent gluon helicity distribution $\Delta g(x)$, more precisely its lowest moment $\Delta G=\int dx\Delta g(x)$. Note that unlike in the case of the quark distribution, the lowest moment of the gluon helicity distribution is still nonlocal. Only in the lightcone gauge does it reduce to a matrix element of the local operator $\vec{E}\times\vec{A}$. Phenomenological global analyses of experimental data indicate that $\Delta G$ is approximately 0.2 at the $\overline{\mathrm{MS}}$ scale $\mu^{2}=10\ \mathrm{GeV}^{2}$, provided that contributions from the small-$x$ region ($x<0.05$) with large uncertainties are excluded de Florian et al. (2009, 2014). On the other hand, recent theoretical developments Ji et al. (2013); Hatta et al. (2014); Ji et al. (2015), particularly within the framework of Large-Momentum Effective Theory (LaMET) Ji (2013, 2014); Ji et al. (2021a), have enabled a direct computation of $\Delta G$ using lattice QCD. In this approach, instead of computing in the IMF and lightcone gauge which are both inaccessible on the lattice, one computes the large-momentum nucleon matrix element of a static gluon spin operator, $\vec{E}\times\vec{A}$, in a physical gauge that leaves the transverse gauge field intact. Actually, there exists an infinite number of operators that can be used to extract $\Delta G$, all belonging to the same universality class Hatta et al. (2014). The static gluon spin matrix element can then be matched to $\Delta G$ through a perturbative factorization or matching relation Ji et al. (2013). Based on this, the $\rm\chi QCD$ collaboration conducted an exploratory calculation of $\Delta G$ by evaluating the nucleon matrix element of $\vec{E}\times\vec{A}$ in the Coulomb gauge, which yields the result $0.251(47)(16)$ Yang et al. (2017). However, the one-loop lattice perturbative renormalization employed in that study introduces a large correction, indicating the necessity of a proper nonperturbative renormalization – which was, at the time, very challenging to implement due to the Coulomb gauge fixing. Furthermore, recent studies Pang et al. (2024) have revealed that the LaMET factorization for $\Delta G$ is incompatible with that for $\Delta g(x)$.
@@
To circumvent the difficulties mentioned above, two new proposals have been made in Ref. Pang et al. (2024). One is based on using the topological current, and the other employs appropriate static nonlocal gluon operators generating the gluon helicity distribution. Both of them have trivial matching only, and therefore resolve the inconsistency previously encountered in the matching for $\Delta G$ and $\Delta g(x)$. Moreover, their nonperturbative renormalization can be easily implemented. This opens a new avenue for systematic studies of $\Delta G$ from lattice QCD.
@@
In this work, we present a state-of-the-art lattice calculation of the total gluon helicity contribution to the proton spin, $\Delta G$, using the topological current approach proposed in Ref. Pang et al. (2024). The calculation is done on ensembles at three different lattice spacings $a=\{0.08,0.09,0.11\}$ fm, which allow us to perform a continuum extrapolation. Distillation Peardon et al. (2009), momentum smearing Bali et al. (2016); Egerer et al. (2021), and cluster decomposition Liu et al. (2018) techniques have been used to improve the signal-to-noise ratio of proton matrix elements and parton Green’s functions, respectively. The nonperturbative renormalization has been performed in the $\mathrm{RI/MOM}$ scheme, following the operator mixing pattern given in Ref. Pang et al. (2024). After converting to the $\overline{\rm MS}$ scheme, the result is then evolved to the scale $\mu^{2}=10\,{\rm GeV}^{2}$ using the renormalization group (RG) equation with three-loop anomalous dimensions.
@@
Theoretical Framework: Following Ref. Pang et al. (2024), the total gluon helicity $\Delta G$ can be extracted by computing the matrix element of the local topological current
@@
in a longitudinally polarized proton. More specifically, the proton matrix element of the Coulomb gauge fixed operator $K^{\mu}$ ($\mu=t,i$) reduces to $\Delta G$ in the IMF, namely,
$\frac{\langle\mathrm{PS}_{\mathrm{Proj.i}}|\int d^{3}xK^{\mu}_{\mathrm{C.G.}}(x)|\mathrm{PS}_{\mathrm{Proj.i}}\rangle}{2S^{\mu}\langle\mathrm{PS}_{\mathrm{Unproj.}}|\mathrm{PS}_{\mathrm{Unproj.}}\rangle}|_{\mathrm{IMF}}=\Delta G,$ (2)
$|\mathrm{PS}_{\mathrm{Unproj.}}\rangle$ denotes the proton external state with a specific momentum but without polarization. In contrast, $|\mathrm{PS}_{\mathrm{Proj.i}}\rangle$ represents the state with specific momentum and a definite polarization direction $i$. The $\mu$ component of the corresponding polarization vector in the $i$th-direction is denoted by $S^{\mu}$. The subscript $\mathrm{C.G.}$ indicates the Coulomb gauge fixing for the topological current.
As shown in Ref. Pang et al. (2024), the ultraviolet (UV) divergence of $K^{\mu}$ is gauge independent. Therefore, we can perform a nonperturbative renormalization of $K^{\mu}$ in the $\mathrm{RI/MOM}$ scheme Martinelli et al. (1995) where the renormalization factor is typically computed in the Landau gauge. $K^{\mu}$ mixes only with the axial-vector current $J^{\mu}_{5}=\bar{q}\gamma_{5}\gamma^{\mu}q$ even under RI/MOM renormalization Pang et al. (2024). The renormalization constants can be determined by the following conditions
$\begin{pmatrix}K^{\mathrm{tree},g/q}\\ J_{5}^{\mathrm{tree},g/q}\end{pmatrix}=\begin{pmatrix}Z_{11}^{\mathrm{RI}}&Z_{12}^{\mathrm{RI}}\\ Z_{21}^{\mathrm{RI}}&Z_{22}^{\mathrm{RI}}\end{pmatrix}\begin{pmatrix}K^{\mathrm{lat.},g/q}\\ J_{5}^{\mathrm{lat.},g/q}\end{pmatrix}$, (3)
where we introduce the abbreviations $\mathcal{O}^{\mathrm{tree/lat.},{g/q}}\equiv\langle{g/q}|\mathcal{O}|{g/q}\rangle^{\mathrm{tree/lat.}}$ to simplify the notation, where $|g\rangle$ or $|q\rangle$ denotes either gluon or quark external states with specific momentum and polarization.
Note that $K^{\mathrm{tree},q}=J_{5}^{\mathrm{tree},g}=0$. Moreover, since only disconnected quark loops are involved, we have $J_{5}^{\mathrm{lat.},g}\approx 0$. Thus, the RI/MOM renormalization constants can be determined as:
@@
where $\mu_{\rm RI}$ denotes the renormalization scale in the RI/MOM scheme. With the above approximation, $Z_{22}^{\mathrm{RI}}$ reduces to the axial-vector current renormalization constant $Z_{A}^{\mathrm{RI}}$ in Ref. Hu and others (2024). The ratio $K^{\mathrm{lat.,q}}/J_{5}^{\mathrm{lat.,q}}$ appearing in the off-diagonal element $Z_{12}^{\mathrm{RI}}$ is of $\mathcal{O}(\alpha_{s})$. Thus, the $\mathrm{RI/MOM}$-renormalized gluon and quark helicities are given by:
@@
The RI/MOM renormalized results can then be converted to the $\overline{\mathrm{MS}}$ scheme via Pang et al. (2024)
@@
where again the scale dependence has been suppressed for simplicity. In the present work, we have included the conversion factors up to three-loops. Their detailed expressions are given in the Supplemental Material A 1.
Lattice Calculation: Our calculation is carried out on the $n_{f}=$2+1 flavor full QCD ensembles generated by CLQCD using the tadpole improved tree level Symanzik (TITLS) gauge action and the tadpole improved tree level Clover (TITLC) fermion action Hu and others (2024); Du and others (2025). The ensembles used for the simulation are described in Table 1. The values of $m_{\pi}$ of all the ensembles are chosen to ensure $m_{\pi}L\geq 3.7$, thereby efficiently suppressing finite-volume effects. For more details, we refer the readers to the Supplemental Material B 1.
@@
Table 1: Lattice setup used for the simulation, with gauge coupling $\hat{\beta}=10/(g_{0}^{2}u_{0}^{4})$, lattice spacing $a$, dimensionless lattice size $\tilde{L}^{3}\times\tilde{T}$, corresponding pion mass $m_{\pi}$, and number of configurations $N_{\mathrm{cfg}}$.
@@
For proton external states at low momenta ($p\leq 0.5\ \mathrm{GeV}$), we implement the distillation smearing technique Peardon et al. (2009), while for higher momentum states we employ an approach combining distillation with momentum smearing Bali et al. (2016); Egerer et al. (2021). We insert the topological current operator into the two-point function of a proton polarized along the $i$th-direction to obtain the disconnected three-point function $C_{3}(t_{f},t_{i},t_{s})$, where $t_{f}$, $t_{i}$, and $t_{s}$ correspond to sink, current insertion, and source time-slices, respectively. The bare matrix elements (BMEs) of the topological current are extracted through the following ratio:
@@
with $C_{2}(t_{f},t_{s})$ representing unpolarized proton two-point correlation functions. We sum over all source time-slices: $R_{K^{\mu}}(t_{f},t_{i})=\sum_{t_{s}}R_{K^{\mu}}(t_{f},t_{i},t_{s})$, and fit the result to the form
$R_{K^{\mu}}(t_{f},t_{i})=\langle K^{\mu}\rangle_{N}^{\mathrm{B.}}+c_{1}e^{-\Delta E(t_{f}-t_{i})}+c_{2}e^{-\Delta Et_{i}},$ (8)
where $\langle K^{\mu}\rangle_{N}^{\mathrm{B.}}$ is defined as nuclear BME divided by the field operator’s normalization factor $2E$, and $c_{1},c_{2},\Delta E$ are fitting parameters. For illustrative purposes, we show in FIG. 1 the fitted ground state BME for both $K_{i}$ and $K_{t}$ at $p=0.98$ GeV. The fitting results for other momenta can be found in the Supplemental Material B 1.
@@
Figure 1: The ratio $R_{K^{\mu}}(t_{f},t_{i})$ at $t_{f}\in[0.63,1.16]\ \mathrm{fm}$ and the fitted BME $\langle K^{\mu}\rangle_{N}^{\mathrm{B.}}$ (the gray error band) for proton external states with p=0.98GeVp=0.98\ $\mathrm${GeV} and ensemble C24P29, utilizing the fitting procedure described in Eq. (8).
@@
$S_{g}^{\mu\nu}(p)=\mathrm{C_{V}}\langle\mathrm{Tr}[A^{\mu}(p)A^{\nu}(-p)]\rangle,$ (10)
@@
where $\mathrm{C_{V}}=2/[(N^{2}_{c}-1)\tilde{V}$], with $N_{c}=3$ representing the number of colors and $\tilde{V}=\tilde{L}^{3}\times\tilde{T}$ denoting the dimensionless lattice volume. In addition, we can define the gluon field renormalization constant
$Z_{g,\mathrm{diag}}^{\mathrm{RI}}(\mu_{\rm RI}^{2})$ $=$ $\frac{\sum_{p_{\rho}=0}1}{\hat{p}^{2}\sum_{p_{\rho}=0}S_{g}^{\rho\rho}(p^{2})},$ (11)
@@
where $\hat{p}_{\mu}=\frac{2}{a}\sin{\frac{p_{\mu}a}{2}}$ represents the lattice momentum, and $\mu_{\rm RI}^{2}=p^{2}$ denotes the RI/MOM renormalization scale. $Z_{11}^{\mathrm{RI}}$ in Eq.(Total Gluon Helicity Contribution to the Proton Spin from Lattice QCD) then takes the following form
$Z_{11}^{\mathrm{RI}}(\mu_{\rm RI}^{2})=\frac{Z_{g,\mathrm{diag}}(\mu_{\rm RI}^{2})}{\mathrm{Tr}[S_{g}^{-1}(p_{1})G_{K}(p_{1},p_{2})S_{g}^{-1}(p_{2})(K^{\mathrm{tree,g}})^{-1}]},$ (12)
@@
where the external momenta are chosen to satisfy $p_{1}=p_{2}=p$. The Green’s function of $K^{\mathrm{tree,g}}$ reads $i\epsilon^{\rho\mu\sigma\nu}\hat{p}_{\sigma}$ in our normalization convention. We hide the $\mu\rho\nu$ indicator of the bare Green’s function $G_{K}$ and bare gluon propagator $S_{g}$ in Eq. (12). $Z_{11}^{\rm RI}$ then reads,
$Z_{11}^{\mathrm{RI}}(\mu_{\rm RI}^{2})=\sum_{\mu,\rho,\nu,\sigma}\frac{i\epsilon^{\rho\mu\sigma\nu}\hat{p}_{\sigma}(S_{g}^{\mu\mu}+S_{g}^{\nu\nu})(p^{2})}{2\mathrm{C_{V}}\hat{p}^{2}G_{K}^{\mu\rho\nu}(p,p)}|_{\mu_{\rm RI}^{2}=p^{2}}.$ (13)
@@
The summation over indices $\mu,\nu,\rho,\sigma$ in conjunction with the Levi-Civita antisymmetric tensor yields a total of 12 distinct configurations. To enhance the signal quality of the pure gluon Green’s function $G_{K}(p,p)$ in Eq. (13), we abandoned original brute force calculation (which requires huge statistics) and instead adopted the Cluster Decomposition Error Reduction (CDER) technique Liu et al. (2018), which requires a large volume ensemble.
Due to the limited volume of C24P29, we have chosen to compute the renormalization constants on a larger volume ensemble, C48P23, which shares the same lattice spacing. In contrast, for the other two ensembles E32P29 and F32P30, the CDER can be directly applied. For more details on the CDER implementation, we refer readers to the Supplemental Material C 1.
@@
The conversion factors from the RI/MOM scheme to the $\overline{\rm MS}$ scheme are evaluated by choosing $\mu/\mu_{\rm RI}=e^{-{5}/{6}}$ Brodsky et al. (1983) in the fixed-order expression up to three-loops, and then RG-evolve to $\mu^{2}=10\,{\rm GeV}^{2}$. Subsequently, the discretization effects of ${\cal O}(a^{2}p^{2})$ are removed following the strategy proposed in Ref. Yang et al. (2018b). Note that here the extrapolation is conducted by varying $p^{2}$ (and thus $a^{2}p^{2}$) and determining the value at $a^{2}p^{2}=0$. To simultaneously ensure the reliability of RI/MOM renormalization and an acceptable signal-to-noise ratio after applying the CDER, we choose the data in the momentum interval $\mu_{\mathrm{RI}}\in[3.0,5.0]\ \mathrm{GeV}$ for the extrapolation, which correspond to $a^{2}p^{2}\in[2.6,7.1]$ for ensemble C48P23 (the CDER cutoff parameters for $K^{\mu}$ and $A^{\mu}$ are 1.26fm1.26\ $\mathrm${fm} and 0.74fm0.74\ $\mathrm${fm}, respectively). The extrapolated result for the inverse of $\bar{Z}_{11}^{\mathrm{lat}}\equiv R_{11}Z_{11}^{\mathrm{RI}}$ is shown in Fig. 2. Similar analyses have also been performed for the other two ensembles with different lattice spacings.
@@
The mixing with quark spin $\Delta\Sigma$, turns out to be an order of magnitude less than $\bar{Z}_{11}^{\mathrm{lat}}$, especially at large $a^{2}p^{2}$ (See the Supplemental Material A 1 for more details). These properties offer a significant simplification in our numerical analysis, leading us to